OpenStudy (anonymous):

What is the solution to the equation? c+(-2)=-0.7 c=?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what did you get?

OpenStudy (anonymous):

well I was thinking maybe try crossing it idk I am so stuck I tried -2 + 0.7 I got -1.3 but I dont think its right lol

OpenStudy (anonymous):

yeah its not lol

jimthompson5910 (jim_thompson5910):

hint: c+(-2) is the same as c-2

jimthompson5910 (jim_thompson5910):

c-2 = -0.7 c = ???

OpenStudy (anonymous):

2.7

OpenStudy (anonymous):

GEEZ THAT TOOK ME FOREVER TO FIGURE OUT! lol

jimthompson5910 (jim_thompson5910):

you add 2 to both sides

jimthompson5910 (jim_thompson5910):

it's not c = 2.7

OpenStudy (anonymous):

-2.7

jimthompson5910 (jim_thompson5910):

c-2 = -0.7 c-2+2 = -0.7+2 c = -0.7+2 c = ???

OpenStudy (anonymous):

1.3

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

lol I am not very good at math as u can tell

jimthompson5910 (jim_thompson5910):

you'll get better with practice